OpenStudy (anonymous):

A solution was made by diluting 80 g of NaOH in a 500cm^3 volumetric flask with distilled water and bringing it up to the mark. Calculate the Mass Concentration. mole concentration, the no. of moles of NaOH present in 25 cm^3 and the concentration of a new solution of NaOH made up by diluting 25 cm^3 of the first solution in a 250 ml volumetric flask. I only need the answers.....i already worked it out...just want to know if my answers are correct! thanks!!

thomaster (thomaster):

All right can you post your work? i'll check if it's correct :)

OpenStudy (abb0t):

Lol. Sneaky snake kym...

OpenStudy (anonymous):

ok mass concentration= 160 gdm^3.......mole concentration=4 mol dm^3.....the # of moles of NaOH=0.1 mol and the concentration of the new solution=0.4 mol dm^3... That's it!

OpenStudy (anonymous):

@thomaster

thomaster (thomaster):

Okay 500cm^3=0.5 liter The w/v concentration is 80/0.5 = 160 g/L so correct NaOH has a Mw of about 40g/mol. You have 160 gram/L,, 160/40 = 4 mol/Liter = 4M so correct Number of moles: 25cm^3=0,025 liter. you have a 4M NaOH concentration so 4 mol/liter 0,025 * 4 = 0.1 so correct. New solution: dilution of 0.025 L in 0.250 Liter = 10x so 4M/10 = 0.4M correct

OpenStudy (anonymous):

oh great! everything's correct! thanks! :)